# MedAgentBench: Dataset for Benchmarking LLMs as Agents in Medical Applications

### v1

@@ Abstract
Methods To address this gap to the deployment of agentic AI in healthcare, we introduce MedAgentBench, a broad evaluation suite designed to assess the agent capabilities of large language models within medical records contexts. MedAgentBench encompasses 100 patient-specific clinically-derived tasks from 10 categories written by human physicians, realistic profiles of 100 patients with over 700,000 data elements, a FHIR-compliant interactive environment, and an accompanying codebase. The environment uses the standard APIs and communication infrastructure used in modern EMR systems, so it can be easily migrated into live EMR systems.
Results MedAgentBench presents an unsaturated agent-oriented benchmark that current state-of-the-art LLMs exhibit some ability to succeed at. The best model (GPT-4o) achieves a success rate of 72%. However, there is still a substantial space for improvement to give the community a next direction to optimize. Furthermore, there is significant variation in performance across task categories.
@@ 1 Introduction
Recent large language models (LLMs) have demonstrated significant advancements, particularly in their ability to serve as agents via active task execution thereby surpassing their traditional role as chatbots [1]. These tasks can be specified at a high level, and agents can dynamically interpret high-level instructions. Agents can then leverage their planning, reasoning, iterative refinement and tool utilization capabilities autonomously to address these tasks. For example, you can ask the agent about the current weather, and the agent might do a web search or call a weather service API and give you an up-to-date answer. Another example is that you can ask the agent to schedule a meeting with certain people, and the agent might send a calendar invite to attendees and modify your calendar to reflect that.
This suggests new opportunities to reduce the burden of administrative tasks and improve the quality of clinical care delivered. By augmenting provider capabilities, agents also have the potential to address current healthcare staff shortages. Examples of potential agentic workflows in healthcare include assessing preoperative risk with regard to surgical candidacy[2], appointment scheduling for patients[3], surveillance of regulatory compliance with hospital safety measures[4], clinical triage[5], electronic health record configuration[6] and insurance prior authorization[7].
@@
In the medical domain, traditional QA-based AI benchmarks such as MedQA, MedMCQA [16] are saturated with impressive performance, and models show superhuman performance on some structured clinical reasoning tasks [17]. Some work like CRAFT-MD [18] and AgentClinic[19] argues that evaluation like structured medical exams is an over-simplification of the real-world interaction between clinicians and patients, and [20] shows LLMs lack metacognition. Salient and difficult benchmarks also help model developers track progress and end users with model selection [21]. Therefore, we need a benchmark for more advanced capabilities, such as agent capabilities in complex environments.
@@
    Dataset We create and release a broad evaluation suite named MedAgentBench, aiming to benchmark large language models for their agent capabilities beyond traditional question answering. It consists of 100 clinically-relevant and verifiable tasks from 10 categories written by licensed human clinicians.
@@
    Benchmark Results We evaluate 9 state-of-the-art large language models (GPT-4o, GPT-4o mini, Gemini 2.0 Flash, Claude 3.5 Sonnet, DeepSeek-V3, Qwen2.5, Llama 3.3, Gemma2 and Mistral v0.3) using MedAgentBench to establish to the current progress. Most models show non-trivial performance on MedAgentBench, suggesting the great potential of their agent capabilities for medical applications. However, they are not yet ready to serve as highly reliable agents. Furthermore, there is significant variation in performance across task categories.
@@ 2 MedAgentBench
Figure 1: Schematic diagram of MedAgentBench architecture.
@@ 2.1 Tasks
Two internal medicine physicians (KB, JHC) submitted 100 clinically derived tasks commonly encountered that could benefit from computer agent automation. Tasks were curated by level of complexity and clinical relevance. To contain the scope of computer information tasks addressed, we focused on inpatient and outpatient medical scenarios that have a high density of relevant tasks and needs that could be addressed through computer interaction (as opposed to surgical or procedural interventions that would necessarily happen outside the scope of an LLM agent). Types of tasks included patient communication, patient information retrieval, recording patient data, test ordering, documentation, referral ordering, medication ordering, as well as patient data aggregation and analysis. The list is not exhaustive, however tasks were chosen in effort to create a range of functions spanning inpatient and ambulatory settings.
Table 1: MedAgentBench task categories and example instructions.
Broad category Example Instruction Example Context
@@
Task structure typically included elements such as patient MRN, timing of request (“over last 24 hours”), and data to be recorded (blood pressure value). We also included NDC, LOINC, base names, and SNOMED codes where applicable.
@@ 2.2.1 Patient cohort
We randomly sample 100 patients from a cohort with an inpatient sodium lab test ordered on the morning of November 13, 2023. The characteristics of the cohort are summarized in Table 2.
Table 2: Characteristics of patient cohort
@@ 2.4.1 Metrics
We use task success rate as the main evaluation metric, as it is commonly used in agent benchmarks. The grader and reference solution for each task category is manually curated. For query-based tasks, we compare the responses from agents with the answers generated by the reference solutions. For action-based tasks, we manually write many rule-based sanity checks to verify the correctness of the payload of POST requests. If the agent system requests for invalid actions or exceeds the maximum number of interaction rounds, it is considered as failed.
@@ 2.4.2 Models
We select a variety of state-of-the-art LLMs across different providers and sizes for benchmarking. They include GPT-4o, GPT-4o mini from OpenAI, Gemini 2.0 Flash from Google, Claude 3.5 Sonnet v2 from Anthropic, DeepSeek-V3 from DeepSeek, Qwen2.5 from Alibaba, Llama 3.3 from Meta, Gemma2 from Google and Mistral v0.3 from Mistral AI (via Together AI serverless API). We set the temperature to zero for all models.
2.4.3 Agent system
We develop a simple agent system to establish the baseline performance, inspired by BFCL [10]. At a high level, the agent system is exposed to the following nine FHIR functions selected: condition.search, lab.search, vital.search, vital.create, medicationrequest.search, medicationrequest.create, procedure.search, procedure.create and patient.search. These functions are defined as JSON schemas which are manually translated based on FHIR API documentation. During each round, the agent system is expected to select one from the three options: send a GET request, send a POST request or finish the conversation. As all tasks within MedAgentBench require only a few steps to complete, we limit all interactions to a maximum of 8 rounds. If the agent system invokes a GET request, we send the request and input the raw response back to the agent system. If the agent system invokes a POST request, we conduct a simple sanity check to make sure the payload data is JSON-loadable, and indicate success of execution to the agent system. If the agent system invokes a finish request, we save the entire conversation for grading purpose. The specific prompt used is included in the appendix. Gemini 2.0 Flash tends to encapsulate the code in a ```tool_code block, so we remove the block separators before parsing.
@@ 2.5 Main results
The performance of five state-of-the-art LLMs on MedAgentBench is shown in Table 3. Most models show non-trivial performance on MedAgentBench, with GPT-4o performing the best with an overall success rate of 76%. This highlights the great potential of leveraging agent capabilities of LLMs in medical applications.
@@
Table 3: Success rate (SR) of state-of-the-art LLMs on MedAgentBench.
@@
GPT-4o N/A API 72% 76% 68%
GPT-4o-mini N/A API 71% 66% 76%
Claude 3.5 Sonnet v2 N/A API 70% 84% 56%
DeepSeek-V3 685B open 56% 60% 52%
Qwen2.5 72B open 53% 44% 62%
Llama 3.3 70B open 49% 54% 44%
Gemma2 27B open 28% 40% 16%
Gemini 2.0 Flash N/A API 19% 32% 6%
Mistral v0.3 7B open 2% 4% 0%
2.6 Analysis
Among the 100 tasks in MedAgentBench, half (50) only require information retrieval via GET requests, while the other half require the modification of medical records through POST requests (often in combination with GET requests beforehand). We calculate task success rates for these two subgroups and name them as query SR and action SR respectively.
Most models, except GPT-4o-mini, are better at query-based tasks than action-based tasks, suggesting that we can start exploring use cases which only require information retrieval first.
Gemini 2.0 Flash only achieves a low success rate in MedAgentBench. One common error pattern of Gemini 2.0 Flash is the model does not follow the instruction exactly, as the agent outputs invalid actions in 36% of the cases. For example, the model tends to output the code in a tool_code or json block, although the instruction has stated that no other text should be in the response. Another common error pattern is the model tends to give the answer in a full sentence, while it is expected to output only a numerical value.
@@ 3 Discussion
MedAgentBench is a benchmark dataset to drive progress in leveraging agent capabilities of large language models for medical applications. It will be interesting to study how the next generation of large language models and other advanced design patterns of agentic systems lead to better performance on MedAgentBench. There is a trade-off between the number of tasks and cost for evaluation. We decided that the first release of MedAgentBench would contain 100 tasks and 100 patient profiles to achieve accurate estimates of performance at reasonable prices.
@@ Appendix A Appendix
A.1 Prompts for the agent system

### v2

# MedAgentBench: A Realistic Virtual EHR Environment to Benchmark Medical LLM Agents

@@ Abstract
Methods To address this gap to the deployment of agentic AI in healthcare, we introduce MedAgentBench, a broad evaluation suite designed to assess the agent capabilities of large language models within medical records contexts. MedAgentBench encompasses 300 patient-specific clinically-derived tasks from 10 categories written by human physicians, realistic profiles of 100 patients with over 700,000 data elements, a FHIR-compliant interactive environment, and an accompanying codebase. The environment uses the standard APIs and communication infrastructure used in modern EMR systems, so it can be easily migrated into live EMR systems.
Results MedAgentBench presents an unsaturated agent-oriented benchmark that current state-of-the-art LLMs exhibit some ability to succeed at. The best model (Claude 3.5 Sonnet v2) achieves a success rate of 69.67%. However, there is still substantial room for improvement which gives the community a next direction to optimize. Furthermore, there is significant variation in performance across task categories.
@@ 1 Introduction
Figure 1: Schematic diagram of MedAgentBench framework. The MedAgentBench workflow begins with a clinician specifying a high-level task, after which the agent orchestrator interacts with both the LLM provider and the electronic medical record environment to finish the task and finally provide feedback to the clinician.
Recent large language models (LLMs) have demonstrated significant advancements, particularly in their ability to serve as agents via active task execution thereby surpassing their traditional role as chatbots [1, 2]. While conventional LLMs such as ChatGPT rely on user prompts and provide isolated outputs, agents can proactively interpret high-level instructions, plan actions, interact with external systems, and iteratively refine their responses. This transition marks a fundamental shift from AI as a tool to AI as a teammate, capable of maintaining memory, integrating contextual knowledge, and orchestrating specialized tools within complex environments [3]. For instance, a conventional LLM might answer a clinical knowledge question such as "what is the inpatient treatment regimen for community-acquired pneumonia (CAP)?" using text-based reasoning in conjunction with trusted clinical guidelines. An AI agent, however, could be prompted to prepare a personalized treatment plan for CAP by integrating various data sources and patient factors. At which point, an agent would calculate personalized patient risk scores, assess for Pseudomonas infection risk factors, analyze for potential medication interactions and allergies, incorporate prior culture data and local antibiogram information, and subsequently queue antibiotics and other supportive care orders for the physician to review and sign. Similarly, an agent can autonomously schedule follow-up visits by integrating with clinical workflows, rather than merely providing a scheduling recommendation [4].
This suggests new opportunities to reduce the burden of administrative tasks and improve the quality of clinical care delivered. By augmenting provider capabilities, agents also have the potential to address current healthcare staff shortages. Examples of potential agentic workflows in healthcare include assessing preoperative risk with regard to surgical candidacy[5], surveillance of regulatory compliance with hospital safety measures[6], clinical triage[7], electronic health record configuration[8] and insurance prior authorization[9].
@@
In the medical domain, traditional QA-based AI benchmarks such as MedQA, MedMCQA [18] are saturated with impressive performance, and models show superhuman performance on some structured clinical reasoning tasks [19]. Some work like CRAFT-MD [20] and AgentClinic[21] argues that evaluation like structured medical exams is an over-simplification of the real-world interaction between clinicians and patients, and [22] shows LLMs lack metacognition. Training datasets for medical contexts have been created to improve performance for complex medical scenarios[23]. Salient and difficult benchmarks also help model developers track progress and end users with model selection [24]. Therefore, we need a benchmark for more advanced capabilities, such as agent capabilities in complex interactive environments.
@@
    Dataset We create and release a broad evaluation suite named MedAgentBench, aiming to benchmark large language models for their agent capabilities beyond traditional question answering. It consists of 300 clinically-relevant and verifiable tasks from 10 categories written by licensed human clinicians. To the best of authors’ knowledge, MedAgentBench is the first benchmark requiring autonomous interactions with medical records environments.
@@
    Benchmark Results We evaluate 12 state-of-the-art large language models (Claude 3.5 Sonnet, o3-mini, GPT-4o, GPT-4o mini, Gemini 2.0 Pro, Gemini 2.0 Flash, Gemini 1.5 Pro, DeepSeek-V3, Qwen2.5, Llama 3.3, Gemma2 and Mistral v0.3) using MedAgentBench to establish to the current progress. Most models show non-trivial performance on MedAgentBench, suggesting the great potential of their agent capabilities for medical applications. However, they are not yet ready to serve as highly reliable agents. Furthermore, there is significant variation in performance across task categories.
@@ 2 MedAgentBench
A typical envisioned workflow (depicted in Figure 1) for the agentic system would be 1) a clinician specifies a high-level task to the agent orchestrator, 2a) the agent interprets the task, and plans function calls, 2b) the agent executes this by sending requests to the FHIR server to modify, for example, the medical records database, and 3) the agent interface (orchestrator) gives an output to the user summarizing the tasks performed.
@@ 2.1 Tasks
Two internal medicine physicians (KB, JHC) submitted 300 clinically derived tasks commonly encountered that could benefit from computer agent automation. Tasks were curated by level of complexity and clinical relevance. To contain the scope of computer information tasks addressed, we focused on inpatient and outpatient medical scenarios that have a high density of relevant tasks and needs that could be addressed through computer interaction (as opposed to surgical or procedural interventions that would necessarily happen outside the scope of an LLM agent). Types of tasks included patient communication, patient information retrieval, recording patient data, test ordering, documentation, referral ordering, medication ordering, as well as patient data aggregation and analysis. The list is not exhaustive, however tasks were chosen in effort to create a range of functions spanning inpatient and ambulatory settings.
Table 1: Broad task categories in MedAgentBench. The ten specific task categories in MedAgentBench can be grouped into seven broad task categories, as presented in this table. Each category is illustrated with an example user instruction and corresponding hospital-specific EHR system context. Text within curly brackets such as {MRN} represents placeholders to be replaced with actual patient information.
Broad category Example user instruction Example context
@@
Task structure typically included elements such as patient MRN, timing of request (“over last 24 hours”), and data to be recorded (blood pressure value). We also included NDC, LOINC, base names, and SNOMED codes where applicable. Of note, ’instructions’ are written by users (e.g. clinicians) and ’context’ is managed by hospital EHR system administrators, given that many hospitals have EHR configurations unique to their environment. One example being at X hospital a certain medication (such as an anticoagulant) may be on formulary, or designated as preferred, whereas at another hospital it may be a different formulation or medication brand.
@@ 2.2.1 Patient cohort
We randomly sample 100 patients from a cohort with an inpatient sodium lab test ordered on the morning of November 13, 2023. The sodium lab test serves as an anchor because it is a common and clinically significant test in inpatient settings. The characteristics of the cohort are summarized in Table 2.
Table 2: Characteristics of patient cohort.
@@ 2.4.1 Metrics
We use task success rate as the main evaluation metric, as it is commonly used in agent benchmarks. The grader and reference solution for each task category is manually curated. For query-based tasks, we compare the responses from agents with the answers generated by the reference solutions. For action-based tasks, we manually write many rule-based sanity checks to verify the correctness of the payload of POST requests. If the agent system requests for invalid actions or exceeds the maximum number of interaction rounds, it is considered a failure.
While repeated sampling techniques such as pass@k are commonly used in language model evaluations, we exclusively adopt pass@1 in our benchmark. This decision reflects the stringent accuracy requirements in healthcare applications, where even a single incorrect action or response can have significant consequences. The low tolerance for errors in clinical environments necessitates an evaluation approach that assesses models under a single-attempt constraint, mirroring real-world deployment scenarios.
@@ 2.4.2 Models
We select a variety of state-of-the-art LLMs across different providers and sizes for benchmarking. They include o3-mini, GPT-4o, GPT-4o mini from OpenAI, Gemini 2.0 Pro, Gemini 2.0 Flash and Gemini 1.5 Pro from Google, Claude 3.5 Sonnet v2 from Anthropic, DeepSeek-V3 from DeepSeek, Qwen2.5 from Alibaba, Llama 3.3 from Meta, Gemma2 from Google and Mistral v0.3 from Mistral AI (via Together AI serverless API). We set the temperature to zero for all models except o3-mini.
2.4.3 Agent orchestrator
We develop a simple agent orchestrator to establish the baseline performance, inspired by BFCL [12]. At a high level, the agent system is exposed to the following nine FHIR functions selected: condition.search, lab.search, vital.search, vital.create, medicationrequest.search, medicationrequest.create, procedure.search, procedure.create and patient.search. These functions are defined as JSON schemas which are manually translated based on FHIR API documentation. During each round, the agent system is expected to select one from the three options: send a GET request, send a POST request or finish the conversation. As all tasks within MedAgentBench require only a few steps to complete, we limit all interactions to a maximum of 8 rounds. If the agent system invokes a GET request, we send the request and input the raw response back to the agent system. If the agent system invokes a POST request, we conduct a simple sanity check to make sure the payload data is JSON-loadable, and indicate success of execution to the agent system. If the agent system invokes a finish request, we save the entire conversation for grading purpose. The specific prompt used is included in the appendix. Gemini models tends to encapsulate the code in a ```tool_code block, so we remove the block separators before parsing.
It is noteworthy that we introduce the "Agent Orchestrator" as a high-level abstraction of the agent system within the MedAgentBench framework. Developers can implement more complex designs, including compound AI systems with hierarchical reasoning, multiple specialized sub-agents, or memory-augmented decision-making. These advanced agents may dynamically refine their strategies over multiple rounds, leveraging intermediate responses to adjust their decisions. Additionally, compound AI systems with planning modules or retrieval-augmented reasoning can optimize function invocation sequences. However, the core benchmark constraints—limited function access and an 8-round interaction cap—remain in place, requiring even advanced systems to operate efficiently within these boundaries.
@@ 2.5 Main results
The performance of 11 state-of-the-art LLMs on MedAgentBench is shown in Table 3. Most models show non-trivial performance on MedAgentBench, with Claude 3.5 Sonnet performing the best with an overall success rate of 69.67%. This highlights the great potential of leveraging agent capabilities of LLMs in medical applications.
@@
Table 3: Success rate (SR) of state-of-the-art LLMs on MedAgentBench. This table presents the performance of various state-of-the-art large language models (LLMs) on MedAgentBench, measured by overall success rate (SR), query SR, and action SR. The best-performing SR values in each column are highlighted in bold.
@@
Claude 3.5 Sonnet v2 N/A API 69.67% 85.33% 54.00%
GPT-4o N/A API 64.00% 72.00% 56.00%
DeepSeek-V3 685B open 62.67% 70.67% 54.67%
Gemini-1.5 Pro N/A API 62.00% 52.67% 71.33%
GPT-4o-mini N/A API 56.33% 59.33% 53.33%
o3-mini N/A API 51.67% 54.67% 48.67%
Qwen2.5 72B open 51.33% 38.67% 64.00%
Llama 3.3 70B open 46.33% 50.00% 42.67%
Gemini 2.0 Flash N/A API 38.33% 34.00% 42.67%
Gemma2 27B open 19.33% 38.67% 0.00%
Gemini 2.0 Pro N/A API 18.00% 25.33% 10.67%
Mistral v0.3 7B open 4.00% 8.00% 0.00%
2.5.1 Subgroup analysis based on task types
Among the 300 tasks in MedAgentBench, half (150) only require information retrieval via GET requests, while the other half require the modification of medical records through POST requests (often in combination with GET requests beforehand). We calculate task success rates for these two subgroups and name them as query SR and action SR respectively.
Most models, except Gemini 1.5 Pro and Qwen2.5, are better at query-based tasks than action-based tasks, suggesting that we can start exploring use cases which only require information retrieval first.
2.5.2 Common error patterns
Figure 2: Example successful trajectory and common error patterns in MedAgentBench. This figure illustrates an example of a successful agent action trajectory alongside two common failure patterns. (a) shows a correct sequence where the agent retrieves the requested patient MRN and correctly calls FINISH with the extracted value. (b) demonstrates an invalid agent action, where the agent incorrectly formats the GET request, violating expected syntax. (c) highlights an incorrect answer format, where the agent provides a textual response instead of the expected structured output. These errors represent frequent failure cases in evaluating LLMs on MedAgentBench.
Figure 2 shows two common error patterns. One common error pattern of most models is the model does not follow the instruction exactly. For example, Gemini 2.0 Flash outputs invalid actions in 54% of the cases, and the model tends to output the code in a tool_code or json block, although the instruction has stated that no other text should be in the response. Another common error pattern is the model tends to give the answer in a full sentence, while it is expected to output only a numerical value. A concrete example is the model outputs "["value": 5.4]" while the expected answer is "[5.4]".
@@ 3 Discussion
MedAgentBench is a benchmark dataset to drive progress in leveraging agent capabilities of large language models for medical applications. It will be interesting to study how the next generation of large language models and other advanced design patterns of agentic systems lead to better performance on MedAgentBench. There is a trade-off between the number of tasks and cost for evaluation. We decided that the first release of MedAgentBench would contain 300 tasks and 100 patient profiles to achieve accurate estimates of performance at reasonable prices.
@@ Appendix A Appendix
A.1 Screenshot of the interactive environment
Figure  3 shows a screenshot of frontend of the FHIR-compliant interactive environment.
@@ A.1 Screenshot of the interactive environment
A.2 Prompts for the agent system
@@ 3. If you have answered all the questions and finished all the requested tasks, you MUST
A.3 Subgroup analysis based on difficulty level
We further break the tasks into three difficulty levels: easy (requires only one step), medium (requires two steps) and hard (requires at least three steps). Table 4 in the appendix shows a breakdown of performance on different difficulty levels.
Table 4: Success rate (SR) of state-of-the-art LLMs on MedAgentBench by difficulty levels. This table presents the success rates of various large language models (LLMs) on MedAgentBench tasks categorized into three difficulty levels: easy (1 step), medium (2 steps), and hard ($\geq$ 3 steps). The highest success rate in each column is highlighted in bold.
Model Size Form Overall SR Easy SR Medium SR Hard SR
Claude 3.5 Sonnet v2 N/A API 69.67% 100.00% 81.67% 23.33%
GPT-4o N/A API 64.00% 86.67% 70.00% 33.33%
DeepSeek-V3 685B open 62.67% 93.33% 68.33% 24.44%
Gemini-1.5 Pro N/A API 62.00% 82.22% 45.83% 63.33%
GPT-4o-mini N/A API 56.33% 91.11% 55.83% 22.22%
o3-mini N/A API 51.67% 67.78% 65.00% 17.78%
Qwen2.5 72B open 51.33% 72.22% 44.17% 40.00%
Llama 3.3 70B open 46.33% 56.67% 38.33% 46.67%
Gemini 2.0 Flash N/A API 38.33% 98.89% 17.50% 5.56%
Gemma2 27B open 19.33% 33.33% 23.33% 0.00%
Gemini 2.0 Pro N/A API 18.00% 27.78% 14.17% 13.33%
Mistral v0.3 7B open 4.00% 13.33% 0.00% 0.00%